# Local Causal Discovery for Structural Evidence of Direct Discrimination

### v1

@@ Abstract
Fairness is a critical objective in policy design and algorithmic decision-making. Identifying the causal pathways of unfairness requires knowledge of the underlying structural causal model, which may be incomplete or unavailable. This limits the practicality of causal fairness analysis in complex or low-knowledge domains. To mitigate this practicality gap, we advocate for developing efficient causal discovery methods for fairness applications. To this end, we introduce local discovery for direct discrimination (LD3): a polynomial-time algorithm that recovers structural evidence of direct discrimination. LD3 performs a linear number of conditional independence tests with respect to variable set size. Moreover, we propose a graphical criterion for identifying the weighted controlled direct effect (CDE), a qualitative measure of direct discrimination. We prove that this criterion is satisfied by the knowledge returned by LD3, increasing the accessibility of the weighted CDE as a causal fairness measure. Taking liver transplant allocation as a case study, we highlight the potential impact of LD3 for modeling fairness in complex decision systems. Results on real-world data demonstrate more plausible causal relations than baselines, which took 197$\times$ to 5870$\times$ longer to execute.
@@ 1 Introduction
Fairness holds fundamental importance in policy design and algorithmic decision-making, especially in high-stakes domains such as healthcare, insurance, and policing [85, 19]. Various criteria have been proposed to measure unfairness with respect to protected or sensitive attributes, such as gender and ethnicity [93]. Legal doctrines generally differentiate between direct discrimination and indirect or spurious forms of unfairness, codifying the notion that the mechanism of unfairness matters [6, 13]. However, fairness criteria based solely on statistical associations cannot disentangle the direct, indirect, and spurious mechanisms of unfairness [39, 50, 69], limiting their informativeness and actionability from a policy intervention standpoint. Consequently, there is a growing emphasis on applying causal reasoning to questions of fairness [39], shifting the focus from associations to interventions and counterfactual outcomes [5].
Causal fairness analysis provides a theoretical framework for disentangling the mechanisms of unfairness using the language of structural causal models (SCMs) [69]. The decomposition of observed parity gaps into direct, indirect, and spurious effects is a centeral objective [67]. As a result, causal fairness analysis can be framed as a special case of mediation analysis [90]. When the complete ground truth SCM is known, the various pathways of unfairness can be read directly from structural equations or the underlying causal graph (Table A.1). With this structural knowledge in hand, points of potential intervention are made clear and causal measures of unfairness are identifiable.
@@
Previous works on effect estimation in causal fairness [98, 67, 69, 68] and mediation analysis [59, 89, 82, 62] generally assume prior knowledge of the true causal graph. In practice, prior structural knowledge is often incomplete, absent, or contentious for complex domains, even among experts [65]. Even with prior knowledge, mediation analysis is challenging territory. Identifiability in causal mediation has been the subject of complicated discussion and debate among theoreticians [62], raising barriers to entry for applied researchers [89]. Consequently, causal fairness analysis is not easily applied to systems with complex causal models, posing a gap between theory and practice.
When the true SCM is unknown, the underlying structure can be learned from observational data through causal discovery [83, 26]. Global discovery recovers the full causal graph or its equivalence class (e.g., [84, 16, 80, 64, 102]), providing mechanistic insights in the natural and social sciences [73, 46]. However, global discovery can pose modeling challenges, including the curse of dimensionality [100], high sample complexity [84, 71, 42], and exponential time complexity in unconstrained search spaces [84, 16, 17, 18]. Reliability is also an open problem, as global methods can sometimes disagree with expert-defined DAGs in complex domains [79, 65]. In the fairness setting, comparisons across global discovery methods can yield conflicting fairness conclusions [7]. Additionally, given an outcome of interest, global discovery is weighed down by a sort of wastefulness: the relations among all observed variables are learned, whether or not they are causally relevant to the outcome [12].
$X$$Y$$\mathbf{C}$$\mathbf{M}$ Standard Fairness Model
$X$$Y$$\mathbf{C}$$\mathbf{M}$ Direct Mechanism
$X$$Y$$\mathbf{C}$$\mathbf{M}$ Indirect Mechanism
$X$$Y$$\mathbf{C}$$\mathbf{M}$ Spurious Mechanism
Figure 1: The Standard Fairness Model (SFM) is a local subgraph around a protected attribute $X$ and outcome $Y$ (Definition 2.2) [69]. Variables that are irrelevant to fairness analysis are abstracted away, leaving only the set of mediators ($\mathbf{M}$) and the set of confounders ($\mathbf{C}$). Causal fairness analysis reveals whether the dotted paths exist. Direct, indirect, and spurious pathways are in solid red.
@@
Local causal discovery [25, 21] aims to mitigate these inefficiencies by only learning the substructures relevant for specific target variables or downstream tasks, such as causal effect estimation [77, 14, 31, 48, 78] or feature selection [95, 96]. As the Standard Fairness Model (SFM; Figure 1) is itself a local subgraph nucleated around an attribute and outcome of interest [69], local causal discovery offers a natural framework for causal fairness analysis. However, task-specific local discovery algorithms are definitionally one-size-does-not-fit-all. Thus, existing methods may not be optimal for fairness tasks, particularly with respect to time efficiency in the nonparametric setting.
@@
To our knowledge, this work presents the first local causal discovery method specifically tailored for causal fairness analysis. We explicitly address the practicality gap in applying causal fairness theory to complex real-world domains with unknown SCMs, while ensuring implementability and state-of-the-art computational efficiency.
@@
    We introduce local discovery for direct discrimination (LD3), [^1] Code is available on GitHub: https://anonymous.4open.science/r/LD3-4440 a time efficient local causal discovery algorithm that assesses the structural direct criterion (SDC) for direct discrimination detection [69]. LD3 performs a linear number of independence tests with respect to variable set size, with asymptotic guarantees on correctness under sufficient assumptions.
@@
    We define a graphical criterion that is sufficient for identifying the weighted controlled direct effect (CDE) [59, 62], a qualitative indicator of direct discrimination [98]. We prove that this criterion is satisfied by the structural knowledge returned by LD3, enhancing the accessibility of this causal fairness measure.
@@
    As a real-world case study of causal fairness analysis for complex domains, we use LD3 to recover evidence of direct sex-based discrimination in liver transplant allocation in the United States. LD3 recovered more plausible causal relations than local and global baselines [38, 31], which performed 42 to 1021$\times$ more tests and took 197 to 5870$\times$ longer to run.
Paper Structure
Section 2 defines essential concepts in causal fairness analysis and direct effect estimation. In Section 3, we introduce LD3 and describe our theoretical and methodological contributions. Section 4 contextualizes our contributions in the broader literature. Section 5 presents numerical validation, and Section 6 highlights a real-world case study on liver transplant allocation.
@@ 2 Preliminaries
Let univariate random variables be denoted by capital letters (e.g., $X$), with specific values in lowercase (e.g., $X=x$). Multivariate random variables or sets are denoted by boldface capital letters (e.g., $\mathbf{X}$), with vector values in bold lowercase (e.g., $\mathbf{X}=\mathbf{x}$). Graphs or function sets are denoted by calligraphic letters (e.g., $\mathcal{F}$). Let $pa(\cdot)$, $an(\cdot)$, $ch(\cdot)$, and $de(\cdot)$ denote the parent, ancestor, child, and descendant sets for a variable in causal graph $\mathcal{G}$, respectively.
@@ 2.1 Causal Fairness Analysis
Causal fairness analysis can be framed in the language of SCMs [69]. Though the theory of SCMs and the potential outcomes framework are logically equivalent [24, 33, 8], the graphical intuition supplied by SCMs is particularly suited for causal fairness analysis.
Definition 2.1 (Structural causal model (SCM), Pearl [59]).
A SCM is a 4-tuple $\langle\mathbf{V},\mathbf{U},\mathcal{F},P(u)\rangle$ where $\mathbf{U}=\{U_{i}\}_{i=1}^{n}$ denotes a set of exogenous variables determined by factors outside of the model, $\mathbf{V}=\{V_{i}\}_{i=1}^{n}$ denotes a set of observed endogenous variables that are determined by variables in $\mathbf{U}\cup\mathbf{V}$, =$\mathcal{F}=\{f_{i}\}_{i=1}^{n}$ denotes a set of structural functions such that $V_{i}=f_{i}(pa(V_{i}),U_{i})$, and $P(u)$ is a distribution over the exogenous variables $\mathbf{U}$.
Every SCM corresponds to a causal graphical model. To facilitate causal fairness analysis, the true causal graph for an SCM can be projected onto the SFM (Figure 1).
Definition 2.2 (Standard Fairness Model (SFM), Plečko & Bareinboim [69]).
Let $\mathcal{G}=(\mathbf{V},\mathbf{E})$ be a causal graph with vertices $\mathbf{V}$ and edges $\mathbf{E}$. Let $\mathcal{G}_{SFM}$ be the projection of $\mathcal{G}$ onto the SFM. This projection is obtained by selecting a protected attribute-outcome pair $\{X,Y\}\subset\mathbf{V}$ and identifying sets $\mathbf{M},\mathbf{C}\subseteq\mathbf{V}\setminus\{X,Y\}$ that satisfy a specific set of relations with respect to $\{X,Y\}$. Thus, $\mathcal{G}_{SFM}$ is the graph induced by $\{X,Y,\mathbf{M},\mathbf{C}\}\subseteq\mathbf{V}$ that encodes the following assumptions:
  • $X$ is a protected attribute and $Y$ is an outcome of interest.
  • $\mathbf{M}$ is a set of mediators with respect to $X$ and $Y$.
  • $\mathbf{C}$ is a set of confounders with respect to $X$ and $Y$, which may share a latent confounder with $X$ (per the bidirected edge in Figure 1).
Note that $\mathbf{C}$ and $\mathbf{M}$ can be of arbitrary dimensionality, including zero. Probing whether the red paths in Figure 1 exist and estimating the associated effect sizes are central tasks in causal fairness analysis.
@@
Multiple criteria have been proposed for evaluating fairness from graphical structures [69]. For the purposes of this work, we are fundamentally interested in detecting graphical signatures of direct discrimination. We say that an SCM is fair or unfair with respect to direct discrimination when the following structural conditions hold.
Definition 2.3 (Structural direct criterion (SDC), Plečko & Bareinboim [69]).
An SCM is fair in terms of direct discrimination if and only if the protected attribute is not a parent of the outcome.
$SDC=\mathbf{1}(X\in pa(Y))=\begin{cases}1\quad\text{if $X$ is a parent of $Y$,}\\ 0\quad\text{if $X$ is not a parent of $Y$.}\end{cases}$ (1)
2.2 Controlled Direct Effect (CDE)
The CDE is a valid test for identifying direct discrimination, as the true value is non-zero if and only if there is a direct path from the protected attribute to the outcome [98]. Let $X=x$ and $X=x^{*}$ be the exposure value corresponding to treatment and no treatment (i.e., the reference value), respectively.
Definition 2.4 (Controlled direct effect (CDE), Vanderweele [89]).
The CDE measures the expected change in outcome as the exposure changes given covariates $\mathbf{S}$, where mediators $\mathbf{M}$ are uniformly fixed to a constant value $\mathbf{m}$ over the population (as simulated by conditioning on $\mathbf{M}$ [61]).
$CDE(\mathbf{m})$ $\coloneqq\mathbb{E}[Y_{x,\mathbf{m}}-Y_{x^{*},\mathbf{m}}]=\sum_{\mathbf{s}}\big{[}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{]}P(\mathbf{s}).$ (2)
Definition 2.5 (Identifiability conditions of the CDE, Vanderweele [89]).
The CDE is identifiable from observational data when, given covariates $\mathbf{S}$, the following holds for all values $X=x,\mathbf{M}=\mathbf{m}$:
@@ 2.2 Controlled Direct Effect (CDE)
    There is no latent confounding of the exposure and outcome given covariates $\mathbf{S}$, i.e., $Y_{x,\mathbf{m}}\perp\!\!\!\perp X|\mathbf{S}$.
@@
    There is no latent confounding of the mediators and outcome given $\{X,\mathbf{S}\}$, i.e., $Y_{x,\mathbf{m}}\perp\!\!\!\perp\mathbf{M}|X,\mathbf{S}$.
Unlike the natural direct effect, CDE identification does not forbid the existence of confounders for $\{\mathbf{M},Y\}$ that are descended from $X$ (Appendix C in Pearl [62]). There exist pathological examples where adjustment for such confounders induces bias, but unbiased CDE estimates can still be obtained for such adjustment sets when given the appropriate estimator [66, 62, 35] (see Appendix B).
When interaction between the mediator and exposure is present, the CDE does not display effect constancy: the CDE will be non-unique, taking on as many values as there are levels of $\mathbf{M}$. Effect constancy is guaranteed for the CDE in linear systems, but not in nonlinear systems [62]. For the nonparametric setting, we avoid assumptions about the absence of interaction yet still wish to obtain a unique CDE. Thus, we define the weighted CDE as the following expectation over $\mathbf{M}$.
Definition 2.6 (The weighted CDE over $\mathbf{M}$).
For all values of covariates $\mathbf{S}$ and mediators $\mathbf{M}$,
$WCDE$ $\coloneqq\sum_{\mathbf{m}}\sum_{\mathbf{s}}\big{[}\big{[}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{]}P(\mathbf{s})\big{]}P(\mathbf{m}).$ (3)
As the weighted CDE is a weighted sum over elements that are well-defined under the identifiability conditions in Definition 2.5, the weighted CDE is identifiable under Definition 2.5 as well.
@@
As shown in [98], if $X$ has no direct path to $Y$ then $CDE(\mathbf{m})=0$. Thus, $CDE(\mathbf{m})\neq 0$ qualitatively implies direct discrimination. As the weighted CDE is a sum, it can equal zero when all elements are zero or when non-zero elements sum to zero. However, it can only be non-zero if at least one $CDE(\mathbf{m})\neq 0$. Thus, we encourage caution in interpreting zero values.
Selecting Mediators and Confounders for Weighted CDE Estimation
As shown in Pearl [59], the subset of mediators that are parents of $Y$ is sufficient for CDE identification. Similarly, it is sufficient to control only for confounders that are adjacent to $X$ or $Y$, as all backdoor paths definitionally enter into $X$ and $Y$. In this work, we opt for identifying confounders that are parents of $Y$. Note that direct effects are identifiable when the exact causal relations among the selected confounders and mediators are not known [82], which is supportive of our local discovery approach.
Definition 2.7 (Valid adjustment for the weighted CDE).
A valid adjustment set for weighted CDE estimation blocks (1) all backdoor paths for $\{X,Y\}$, (2) all backdoor paths for $\{\mathbf{M},Y\}$, and (3) all mediator paths for $\{X,Y\}$. This blocks non-causal association through confounding paths and indirect causal association through mediator paths, isolating the direct causal effect.

### v3

@@ Abstract
Identifying the causal pathways of unfairness is a critical objective for improving policy design and algorithmic decision-making. Prior work in causal fairness analysis often requires knowledge of the causal graph, hindering practical applications in complex or low-knowledge domains. Moreover, global discovery methods that learn causal structure from data can display unstable performance on finite samples, preventing robust fairness conclusions. To mitigate these challenges, we introduce local discovery for direct discrimination (LD3): a method that uncovers structural evidence of direct unfairness by identifying the causal parents of an outcome variable. LD3 performs a linear number of conditional independence tests relative to variable set size, and allows for latent confounding under the sufficient condition that all parents of the outcome are observed. We show that LD3 returns a valid adjustment set (VAS) under a new graphical criterion for the weighted controlled direct effect, a qualitative indicator of direct discrimination. LD3 limits unnecessary adjustment, providing interpretable VAS for assessing unfairness. We use LD3 to analyze causal fairness in two complex decision systems: criminal recidivism prediction and liver transplant allocation. LD3 was more time-efficient and returned more plausible results on real-world data than baselines, which took 46$\times$ to 5870$\times$ longer to execute.
@@ 1 Introduction
Fairness holds fundamental importance in policy design and algorithmic decision-making, especially in high-stakes domains such as healthcare and policing (Starke et al. 2022; Corbett-Davies et al. 2023). Various criteria have been proposed for measuring unfairness with respect to protected or sensitive attributes, such as gender and ethnicity (Verma and Rubin 2018). Legal doctrines generally differentiate between direct discrimination and indirect or spurious forms of unfairness, codifying the notion that the mechanism of unfairness matters (Barocas and Selbst 2016; Carey and Wu 2022). However, fairness criteria based solely on statistical associations cannot disentangle these mechanisms (Kilbertus et al. 2017; Makhlouf, Zhioua, and Palamidessi 2020), limiting their informativeness and actionability for policy interventions. Consequently, there is a growing emphasis on applying causal reasoning in fairness analysis (Kilbertus et al. 2017), shifting the focus from associations to interventions and counterfactual outcomes.
@@
Causal fairness analysis (CFA) provides a theoretical framework for disentangling the mechanisms of unfairness using the language of structural causal models (SCMs). Previous works in CFA (Zhang and Bareinboim 2018; Plecko and Bareinboim 2023) and the allied field of mediation analysis (Pearl 2014; VanderWeele 2016) generally assume significant prior structural knowledge in order to decompose direct, indirect, and spurious effects. In practice, structural knowledge is often incomplete, absent, or contentious for complex domains, even among experts (Petersen et al. 2023). Furthermore, the identifiability of direct and indirect effects has been a topic of extensive debate among theoreticians (Pearl 2014), raising barriers to entry for applied researchers (Vanderweele 2011). Thus, existing methodologies in CFA can be challenging to apply in complex systems.
Among the many fairness measures proposed in CFA, the controlled direct effect (CDE) is a relatively straightforward qualitative indicator of direct discrimination that takes non-zero values only when the exposure is a direct cause of the outcome (Zhang and Bareinboim 2018). The CDE has often been favored in policy evaluation over alternative direct effect measures (Vanderweele 2011; VanderWeele 2013) as it is more interpretable for real-world interventions and requires fewer untestable assumptions and less prior structural knowledge (Pearl 2001; Shpitser and VanderWeele 2011). To increase the practicality of CFA in complex domains, we choose to focus on the CDE as a starting point for this work.
$X$$Y$$\mathbf{C}$$\mathbf{M}$Direct
$X$$Y$$\mathbf{C}$$\mathbf{M}$Indirect
$X$$Y$$\mathbf{C}$$\mathbf{M}$Spurious
Figure 1: The standard fairness model (SFM) is compactly represented as a local subgraph around protected attribute $X$ and outcome $Y$ (Plečko and Bareinboim 2024). Variables that are irrelevant to CFA are abstracted away, leaving confounders ($\mathbf{C}$) and mediators ($\mathbf{M}$). Directed edges represent the existence of active directed paths, and bidirected edges denote potential confounding. This work aims to identify direct mechanisms of unfairness in a data-driven way.
In low-knowledge domains, we can support CFA by learning causal structure directly from observational data. However, global causal discovery is challenging in finite data due to high sample complexity (Spirtes, Glymour, and Scheines 2001) and exponential time complexity in unconstrained search spaces (Chickering, Heckerman, and Meek 2004; Claassen, Mooij, and Heskes 2013). Learned causal graphs often disagree with expert knowledge in complex domains (Shen et al. 2020; Petersen et al. 2023), and can yield conflicting causal fairness conclusions in CFA (Binkytė et al. 2023 and Section 6 of this paper).
@@
While global discovery learns the relations among all observed variables, local causal discovery only learns the substructures relevant for downstream tasks, such as causal effect estimation (Gupta, Childers, and Lipton 2023; Maasch et al. 2024; Shah, Shanmugam, and Kocaoglu 2024) or feature selection (Yu et al. 2021; Yu, Liu, and Li 2021). As the standard fairness model (SFM; Figure 1) is represented as a local subgraph (Plečko and Bareinboim 2024), local discovery offers a natural framework for CFA. However, since task-specific local discovery algorithms are definitionally one-size-does-not-fit-all, existing methods may not be optimal for fairness tasks.
@@
This work aims to increase the practicality of CFA for direct discrimination in complex domains with unknown causal graphs. Our contributions are three-fold.
@@
    Local discovery for direct discrimination (LD3). This local causal discovery method leverages the problem structure in CFA to efficiently detect graphical signatures of direct discrimination. [^1] Code on GitHub: https://github.com/jmaasch/LD3 LD3 discovers the parents of an outcome variable in a linear number of conditional independence tests with respect to variable set size.
@@
    A graphical criterion for the weighted controlled direct effect (WCDE). This criterion is sufficient to identify a valid adjustment set (VAS) for the WCDE, a qualitative indicator of direct discrimination. This criterion is satisfied by the knowledge returned by LD3.
@@
    Real-world fairness analysis. We deploy LD3 for two fairness problems: (1) racial discrimination in recidivism prediction and (2) sex-based discrimination in liver transplant allocation. LD3 recovered more plausible causal relations than local and global baselines, which performed 11–1021$\times$ more tests and took 46–5870$\times$ longer to run.
@@ 2 Preliminaries
Let capital letters denote univariate random variables (e.g., $X$), with their values in lowercase (e.g., $X=x$). Multivariate random variables or sets are denoted by boldface capital letters (e.g., $\mathbf{X}$), with vector values in bold lowercase (e.g., $\mathbf{X}=\mathbf{x}$). Graphs or function sets are denoted by calligraphic letters (e.g., $\mathcal{F}$). Let $pa(\cdot)$ and $de(\cdot)$ denote the parent and descendant sets for a variable in causal graph $\mathcal{G}$, respectively.
@@ 2.1 Causal Fairness Analysis
CFA can be framed in the language of SCMs and their graphical representations (Plečko and Bareinboim 2024).
Definition 1 (Structural causal model, Bareinboim et al. 2022).
An SCM is a 4-tuple $\langle\mathbf{V},\mathbf{U},\mathcal{F},p(\mathbf{u})\rangle$ where $\mathbf{U}=\{U_{i}\}_{i=1}^{n}$ denotes a set of exogenous variables determined by factors external to the model, $\mathbf{V}=\{V_{i}\}_{i=1}^{n}$ denotes a set of observed endogenous variables determined by $\mathbf{U}\cup\mathbf{V}$, =$\mathcal{F}=\{f_{i}\}_{i=1}^{n}$ denotes a set of structural functions such that $V_{i}=f_{i}(pa(V_{i}),U_{i})$, and $p(\mathbf{u})$ is the distribution over $\mathbf{U}$.
Every SCM corresponds to a causal graphical model. To facilitate CFA, the true causal graph for an SCM can be compactly represented using the SFM (Figure 1).
Definition 2 (Standard fairness model, Plečko and Bareinboim 2024).
Let $\mathcal{G}=(\mathbf{V},\mathbf{E})$ be a causal graph with vertices $\mathbf{V}$ and edges $\mathbf{E}$. Let $\mathcal{G}_{\mathrm{SFM}}$ be the projection of $\mathcal{G}$ onto the SFM, which is obtained by (1) selecting a protected attribute-outcome pair $\{X,Y\}\subset\mathbf{V}$ and (2) identifying sets $\mathbf{M},\mathbf{C}\subseteq\mathbf{V}\setminus\{X,Y\}$ that meet the following conditions:
  • $\mathbf{M}$ is the set of mediators with respect to $X$ and $Y$;
  • $\mathbf{C}$ is the set of confounders with respect to $X$ and $Y$.
Note that $\mathbf{C}$ or $\mathbf{M}$ can be the empty set, and confounding can exist among $\mathbf{C}$, $\mathbf{M}$, $X$, $Y$ (Figure 1).
@@
Multiple criteria have been proposed for evaluating fairness from graphical structures. Here, we focus on a criterion for direct discrimination.
Definition 3 (Structural direct criterion (SDC), Plečko and Bareinboim 2024).
An SCM is fair with respect to direct discrimination if and only if the SDC evaluates to 0:
$SDC$ =$=\begin{cases}1\quad\text{if $X$ is a parent of $Y$,}\\ 0\quad\text{if $X$ is not a parent of $Y$.}\end{cases}$ (1)
2.2 Controlled Direct Effect
The CDE can be used to test for direct discrimination, as the true value is non-zero if and only if there is a direct path from the protected attribute to the outcome (Zhang and Bareinboim 2018). Let $X=x$ and $X=x^{*}$ be the exposure values corresponding to treatment and no treatment, respectively.
Definition 4 (CDE, Pearl 2014).
The CDE measures the expected change in outcome as the exposure changes when mediators $\mathbf{M}$ are uniformly fixed to a constant value $\mathbf{m}$:
$\mathrm{CDE}(\mathbf{m})$ $\coloneqq\mathbb{E}[Y|do(x,\mathbf{m})]-\mathbb{E}[Y|do(x^{*},\mathbf{m})].$ (2)
Definition 5 (Identifiability conditions of the CDE, Vanderweele 2011).
When we have access to a covariate set $\mathbf{S}$ that controls for confounding of both $\{X,Y\}$ and $\{\mathbf{M},Y\}$, the CDE is identifiable under the following conditions. Let $Y_{\check{x},\check{\mathbf{m}}}$ denote the value of $Y$ under $X=\check{x},\mathbf{M}=\check{\mathbf{m}}$.
@@ 2.2 Controlled Direct Effect
    There is no latent confounding of the exposure and outcome given $\mathbf{S}$, i.e., $Y_{\check{x},\check{\mathbf{m}}}\perp\!\!\!\perp X|\mathbf{S}$ for all $\check{x},\check{\mathbf{m}}$.
@@
    There is no latent confounding of the mediators and outcome given $\{X,\mathbf{S}\}$, i.e., $Y_{\check{x},\check{\mathbf{m}}}\perp\!\!\!\perp\mathbf{M}|X,\mathbf{S}$ for all $\check{x},\check{\mathbf{m}}$.
Then, we obtain $CDE(\mathbf{m})$ by [^2] When $\mathbf{S}$ is not sufficient for valid control of confounding for both $\{X,Y\}$ and $\{\mathbf{M},Y\}$, alternative formulae may be required (Vanderweele 2011; Pearl 2014).
$\sum_{\mathbf{s}}\big{(}\mathbb{E}[Y\mid x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y\mid x^{*},\mathbf{s},\mathbf{m}]\big{)}P(\mathbf{s}).$ (3)
Note that the subset of mediators that are parents of $Y$ is sufficient for CDE identification (Pearl 2001).
Comparison to Alternative Measures
Several estimands can capture direct effects, including CDE (Pearl 2001), natural direct effect (NDE; Pearl 2001), and counterfactual direct effect (Ctf-DE; Zhang and Bareinboim 2018). While CDE is an interventional quantity, NDE and Ctf-DE are counterfactual quantities. In this work, we favor CDE as it requires fewer untestable assumptions over the data generating process (Shpitser and VanderWeele 2011) and less structural knowledge than NDE and Ctf-DE (Pearl 2001; Vanderweele 2011; Zhang and Bareinboim 2018). The main objective of this work is to assess whether the protected attribute is a direct cause of the outcome. The CDE, NDE, and Ctf-DE are all non-zero if and only if $X$ directly points toward $Y$. Thus, we advocate for this simpler estimand for practicality. See Appendix A for an extended comparison of estimands.
Remark 1 (Assumptions on Confounding).
Unlike the NDE, CDE identification does not forbid the existence of confounders for $\{\mathbf{M},Y\}$ that are descended from $X$ (Pearl 2014). As done previously (Vanderweele and Vansteelandt 2009), this work assumes structures in which adjusting for confounders of $\{\mathbf{M},Y\}$ does not induce post-treatment bias, allowing us to identify the CDE with the expression provided in Equation 3. If this assumption does not hold, unbiased CDE estimates can still be obtained given alternative estimators (Petersen, Sinisi, and Van Der Laan 2006).
Weighted CDE
When interaction between the mediator and exposure exists, the value of CDE may vary across different mediator values (Pearl 2014). To avoid assumptions about interaction while still obtaining a unique estimate, we define the weighted CDE (WCDE) as the following expectation over $\mathbf{M}$. For simplicity, we assume discrete $\mathbf{M}$ [^3] Definition 6 generalizes to continuous $\mathbf{M}$ by integrating over its possible values. We note that LD3 operates independently of the variable type of $\mathbf{M}$.
Definition 6 (WCDE, Pearl 2000444Though unnamed, this expression is noted on page 131.).
We define WCDE as
$\sum_{\mathbf{m}^{\prime}}\big{(}\mathbb{E}[Y\mid do(x,\mathbf{m}^{\prime})]-\mathbb{E}[Y\mid do(x^{*},\mathbf{m}^{\prime})]\big{)}P(\mathbf{m}^{\prime}),$ (4)
where $\mathbf{M}^{\prime}\subseteq\mathbf{M}$ are parents of $Y$. Per Equation 3, WCDE is identifiable as
$\sum_{\mathbf{m}^{\prime}}\sum_{\mathbf{s}}\big{(}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}^{\prime}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}^{\prime}]\big{)}P(\mathbf{s})P(\mathbf{m}^{\prime}).$ (5)
Definition 7 (VAS for the WCDE).
Given Definitions 5 and 6, a VAS for WCDE estimation blocks (1) all backdoor paths for $\{X,Y\}$, (2) all backdoor paths for $\{\mathbf{M},Y\}$, and (3) all mediator paths for $\{X,Y\}$.
@@
Similar to CDE, WCDE indicates direct discrimination when its value is non-zero. However, a zero value does not guarantee the absence of direct discrimination, as different CDE values may cancel each other out. Thus, we encourage caution when interpreting zero values.
2.3 Mapping Causal Partitions to the SFM
The methods introduced in this work leverage the causal partition taxonomy defined in Maasch et al. (2024) (Table 1). Given an exposure-outcome pair $\{X,Y\}$, any arbitrary variable set $\mathbf{Z}$ can be uniquely partitioned into eight disjoint subsets (which may be empty) that are defined by the types of causal paths that they share with $X$ and $Y$. By focusing structure learning on relationships that are causally relevant to the exposure and outcome, this partition taxonomy provides practical building blocks for local discovery. Mapping this partition taxonomy to the SFM, $\mathbf{Z}_{1}$ are confounders $\mathbf{C}$ (and their proxies) and $\mathbf{Z}_{3}$ are mediators $\mathbf{M}$ (and their proxies). [^5] See Maasch et al. (2024) for more formal partition definitions. Proxy variables are not relevant in this setting, as these cannot be parents of $Y$ and are not returned by LD3. When referring to the union of multiple partitions, we use notation of the form $\mathbf{Z}_{1,3}\coloneqq\mathbf{Z}_{1}\cup\mathbf{Z}_{3}$.